# A Comprehensive Survey on Continual Learning in Generative Models

### v3

@@ Abstract
The rapid advancement of generative models has enabled modern AI systems to comprehend and produce highly sophisticated content, even achieving human-level performance in specific domains. However, these models remain fundamentally constrained by catastrophic forgetting — a persistent challenge where adapting to new tasks typically leads to significant degradation in performance on previously learned tasks. To address this practical limitation, numerous approaches have been proposed to enhance the adaptability and scalability of generative models in real-world applications. In this work, we present a comprehensive survey of continual learning methods for mainstream generative models, encompassing large language models, multimodal large language models, vision-language action models, and diffusion models. Drawing inspiration from the memory mechanisms of the human brain, we systematically categorize these approaches into three paradigms: architecture-based, regularization-based, and replay-based methods, while elucidating their underlying methodologies and motivations. We further analyze continual learning setups for different generative models, including training objectives, benchmarks, and core backbones, offering deeper insights into the field. The project page of this paper is available at https://github.com/Ghy0501/Awesome-Continual-Learning-in-Generative-Models.
Index Terms:
@@
Continual Learning, Generative Models, LLM, MLLM, VLA, Diffusion Models, Catastrophic Forgetting
1 Introduction
@@
The evolution of artificial intelligence is undergoing a paradigm shift from “understanding the world” to “creating the world”. Traditional discriminative models, such as classification networks [1, 2, 3, 4, 5, 6] and object detectors [7, 8, 9, 10, 11], have achieved remarkable success over the past decade. These models are effective at learning decision boundaries from large-scale labeled datasets, enabling accurate recognition and judgment of known concepts. However, this intelligence, based on discrimination, has shown limitations in creativity and generalization. It excels at distinguishing known concepts within a single modality but struggles both to generate novel content and to establish semantic associations across multiple modalities. The rise of generative models addresses these limitations by employing large-scale pre-training to capture the underlying data distribution. As a result, mainstream generative models such as Large Language Models (LLMs) [12, 13, 14], Multimodal Large Language Models (MLLMs) [15, 16, 17, 18, 19], Vision-Language Action (VLA) [20, 21], and Diffusion Models [22, 23] are now capable of not only interpreting input features but also actively generating coherent text, realistic images, and even cross-modal content. As the field transitions from discriminative to generative paradigms, AI is evolving from merely recognizing the known to actively creating the new. This shift is redefining the foundations of human-machine collaboration by enabling more flexible, and multimodal generation capabilities.
Figure 1: (a) Continual learning lets humans learn lifelong without forgetting; AI models aim to do the same while retaining old knowledge. (b) Discriminative models focus on perceptual tasks such as classification, detection, and segmentation, among others, whereas generative models additionally demand the capability to produce novel content grounded in perceptual understanding.
@@
The success of generative models like GPT-4o fundamentally relies on large-scale pre-training and human alignment. However, this process is inherently dynamic, as languages, data distributions, and user demands continually evolve. More importantly, it is infeasible to account for all possible future scenarios in advance, which underscores the need for models to possess continual learning capabilities comparable to those of humans, as illustrated in Figure 1 (a). Such capabilities are therefore essential for generative models to move beyond the closed-world assumption [24] and enable deployment in open-ended, real-world scenarios [25, 26, 27]. A central challenge in this context is catastrophic forgetting, where learning new information causes previously acquired knowledge to be overwritten [28, 29, 30]. For instance, fine-tuning general-domain models (e.g., ChatGPT) on specialized domains (e.g., mathematics) may improve in-domain performance at the expense of forgetting prior capabilities. When models are trained sequentially on multiple downstream tasks, each new update compounds this effect, leading to a progressive degradation in knowledge retention. A straightforward solution is to retain data from previous tasks and perform joint training with new data. However, this approach is often impractical due to significant storage costs [31] and data privacy concerns [32]. In contrast, both children and adults rarely experience catastrophic forgetting when acquiring new knowledge [33, 34], highlighting a fundamental gap between artificial and biological learning systems. To bridge this gap, the study of continual learning [35, 36, 37, 38, 39] has gained increasing attention, aiming to equip models with the ability to acquire knowledge sequentially while preserving previously learned skills.
@@
Compared to traditional discriminative models that focused on perceptual tasks, generative models in continual learning settings are expected to go a step further, as shown in Figure 1 (b), by generating new content based on understanding and recognition, owing to the diversity of modeling objectives and task formats. On the one hand, discriminative methods focus on modeling conditional probabilities (i.e. $P(y|x)$), where catastrophic forgetting primarily manifests as a drift in the decision boundary within the classification space [40, 41, 42, 43, 44]. In contrast, generative models aim to learn joint probability distributions (i.e. $P(x,y)$) or perform autoregressive generation, which requires continual learning to maintain both semantic coherence and knowledge completeness within an open-ended generative space. To achieve this, it is essential to refine the generative reasoning process alongside the underlying factual knowledge representation, thereby mitigating hallucinations [45, 46] and semantic inconsistencies [47]. On the other hand, generative models encounter highly diverse tasks in the continual learning process, with differences arising from factors such as new modality inputs, novel domain knowledge, and other evolving task characteristics. This variability increases the risk of catastrophic forgetting and places greater demands on multi-task knowledge integration. In contrast, discriminative models typically focus on a narrow range of tasks, such as classification, detection, or segmentation, which makes continual learning comparatively less challenging. Figure 2 demonstrates the continual learning scenarios across four representative generative models. Despite their impressive capabilities, enabling these models to learn continuously remains an open and pressing research challenge.
@@
Figure 2: Illustrations of continual learning in generative models: Adapting to real-world demands by learning new instructions, knowledge, modalities, actions, and creations without catastrophic forgetting.
Existing continual learning methods have been extensively developed for discriminative models [48, 49, 50, 51, 52]. In contrast, research on continual learning in generative models has generally focused on specific architectures, such as Large Language Models [53, 54, 55, 56] or Multimodal Large Language Models [57, 58, 59]. Although these works provide systematic overviews of continual learning techniques within their respective domains, they often lack broader discussions and cross-model connections. Motivated by this gap, and considering the growing interest in generative and large-scale unified models within the AI community, we believe that an up-to-date and comprehensive survey will offer a holistic perspective to guide future work. In this paper, we present the first comprehensive review of continual learning in mainstream generative models, systematically analyzing key approaches and challenges across Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models, as illustrated in Figure 2. In addition, we provide a unified perspective by identifying shared methodological dimensions of these generative models in the context of continual learning.
@@
The paper is organized as follows: In Section 2, we present the setup of continual learning, including its basic formulation, evaluation metrics, and taxonomy. In Section 3 to Section 6, we provide an in-depth discussion of continual learning in Large Language Models (Section 3), Multimodal Large Language Models (Section 4), Vision-Language Action Models (Section 5), and Diffusion Models (Section 6). Each domain is discussed in terms of the training strategies for continual learning tasks, common benchmarks, model architectures, as well as the motivations and implementations of existing methods. In Section 7, we further explore future research avenues and emerging trends in continual learning methods based on generative models. Finally, Section 8 concludes the paper and introduces a continuously updated GitHub repository for tracking recent progress in continual learning for generative models.
2 Overview of Continual Learning
@@
(§ 3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=3.0cm [Architecture-based
@@
(§ 3.2), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [TPEM [60], Zheng et al [61], TGL [62], LOIRE [63], AdapterCL [64], SAPT [65], MoRAL [66], I-LoRA [67], SLIM [68],
@@
P-Prompts [69], Q-Tuning [70], WAVE-CRE [71] , color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
@@
(§ 3.3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [O-LoRA [72], DAS [73], ARPER [74], SEEKR [75], TaSL [76],
@@
DYNAINST [77], DaCL [78], FVG [79], color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] [Replay-based
@@
(§ 3.4), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [CT0 [80], Huang et al. [81], Anh et al. [82], InsCL [83], HESIT [84],
@@
Sun et al. [85], D-CPT Law [86], LAMOL [87], HMI-LAMOL [88], PCLL [89], DCL [90], SSR [91], color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] ] [Multimodal Large Language Models
@@
(§ 4), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=3.0cm [Architecture-based
@@
(§ 4.2), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [Eproj [92], TAM-CL [93], MoELoRA [94], Continual LLaVA [95], HiDe-LLaVA [96], LLaCA [97], SMoLoRA [98], CL-MoE [99],
@@
LLaVA-CMoE [100], DISCO [101], TRIPLET [102], ModalPrompt [103],
@@
Fwd-Prompt [104], ColPro [105], CluMo [106], MR-LoRA [107], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
@@
(§ 4.3), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [SCD [108], CS-VQLA [109], Chen et al. [110], CrossSDC [111],
MAFED [112], MoInCL [113], QUAD [114], Model Tailor [115],
@@
SPIDER [116], SEFE [117], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] [Replay-based
@@
(§ 4.4), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [VQACL [118], SGP [119], ProtoGroup [120], Lin et al. [121],
@@
Adapt-$\infty$ [122], OASIS [123], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] ] [Vision-Language Action Models
@@
(§ 5), color=cyan!100, fill=cyan!15, very thick, text=black, text width=3.0cm [Architecture-based
@@
(§ 5.2), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [LIBERO [124], Lotus [125], QueST [126], LEGION [127], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 5.3), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [CAMA [128], M2Distill [129], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] [Replay-based
(§ 5.4), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [iManip [130], RWLA [131], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] ] [Diffusion Models
(§ 6), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=3.0cm [Architecture-based
@@
(§ 6.2), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [C-LoRA [132], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 6.3), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [MuseumMaker [133], LFS-Diffusion [134], CIDM [135], ConceptGuard [136], Jha et al [137], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] [Replay-based

### v4

# Continual Learning for Generative AI: From LLMs to MLLMs and Beyond

@@ Abstract.
The rapid advancement of generative models has empowered modern AI systems to comprehend and produce highly sophisticated content, even achieving human-level performance in specific domains. However, these models are fundamentally constrained by catastrophic forgetting, i.e., a persistent challenge where models experience performance degradation on previously learned tasks when adapting to new tasks. To address this practical limitation, numerous approaches have been proposed to enhance the adaptability and scalability of generative AI in real-world applications. In this work, we present a comprehensive survey of continual learning methods for mainstream generative AI models, encompassing large language models, multimodal large language models, vision-language-action models, and diffusion models. Drawing inspiration from the memory mechanisms of the human brain, we systematically categorize these approaches into three paradigms: architecture-based, regularization-based, and replay-based methods, while elucidating their underlying methodologies and motivations. We further analyze continual learning setups for different generative models, including training objectives, benchmarks, and core backbones, thereby providing deeper insights into the field. The project page of this paper is available at https://github.com/Ghy0501/Awesome-Continual-Learning-in-Generative-Models.
@@
Continual Learning, Generative Models, Large Language Models, Multimodal Large Language Models, Vision-Language-Action Models, Diffusion Models, Catastrophic Forgetting
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
The evolution of artificial intelligence is undergoing a paradigm shift from “understanding the world” to “creating the world”. Traditional discriminative models, such as classification networks (Krizhevsky et al., 2012; Simonyan and Zisserman, 2014; He et al., 2016; Vaswani et al., 2017; Zeng and Yu, 2024; Dosovitskiy et al., 2020) and object detectors (Girshick et al., 2014; Ren et al., 2015; He et al., 2017; Lin et al., 2017; Redmon et al., 2016), have achieved remarkable success over the past decade. These models are competent in learning decision boundaries from large-scale labeled datasets, enabling accurate recognition and judgment of known concepts. However, this intelligence, based on discrimination, has shown limitations in creativity and generation. It excels at distinguishing known concepts within a single modality but struggles both to generate novel content and to establish semantic associations across multiple modalities. The rise of generative AI addresses these limitations by employing large-scale pre-training to capture the underlying data distribution. As a result, mainstream generative AI models such as Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2024), Multimodal Large Language Models (MLLMs) (Liu et al., 2023a; Chen et al., 2023b; Bai et al., 2023a; Liu et al., 2024a; Bai et al., 2025), Vision-Language-Action (VLA) Models (Kim et al., 2024a; Zhou et al., 2025), and Diffusion Models (Rombach et al., 2022; Ho et al., 2020) are now capable of not only interpreting input features but also actively generating coherent text, realistic images, and even cross-modal content. As the field transitions from discriminative to generative paradigms, AI is evolving from merely recognizing the known to actively creating the new. This shift is redefining the foundations of human-machine collaboration by enabling more flexible and multimodal generation capabilities.
@@
The success of generative models like GPT-4 (Achiam et al., 2023) fundamentally relies on large-scale generative pre-training and human alignment. However, this process is inherently dynamic, as languages, data distributions, and user demands continually evolve. More importantly, it is infeasible to account for all possible future scenarios in advance, which underscores the need for models to possess continual learning capabilities comparable to those of humans, as illustrated in Figure 1 (a). Such capabilities are therefore essential for generative models to move beyond the closed-world assumption (Zhang et al., 2020) and enable deployment in open-ended, real-world scenarios (Zhou, 2022; Parmar et al., 2023; Zhu et al., 2024b). A central challenge in this context is catastrophic forgetting, where learning new information causes previously acquired knowledge to be overwritten (McCloskey and Cohen, 1989; McClelland et al., 1995; Kirkpatrick et al., 2017). For instance, fine-tuning general-domain models (e.g., ChatGPT) on specialized domains (like mathematics) may improve in-domain performance at the expense of forgetting prior capabilities. When models are trained sequentially on multiple downstream tasks, each new update compounds this effect, leading to a progressive degradation in knowledge retention. A straightforward solution is to retain data from previous tasks and perform joint training with new data. However, this approach is often impractical due to significant storage costs (Krempl et al., 2014) and data privacy concerns (De Lange et al., 2021). By contrast, children and adults typically do not exhibit catastrophic forgetting when learning new information, even in the absence of explicit recall of prior knowledge (Xue et al., 2010; Chang et al., 2025), highlighting a fundamental gap between artificial and biological learning systems. To bridge this gap, the study of continual learning (Van de Ven et al., 2022; Li and Hoiem, 2017; Lopez-Paz and Ranzato, 2017; Rolnick et al., 2019; Zhu et al., 2021b) has gained momentum, aiming to equip models with the ability to learn new knowledge over time while retaining prior skills, especially in the practical cases where the data from old tasks are limited or even inaccessible, particularly in practical scenarios where data from earlier tasks is limited or entirely unavailable.
Figure 1. (a) Continual learning lets humans learn lifelong without forgetting; AI models aim to do the same while retaining old knowledge. (b) Discriminative models focus on perceptual tasks such as classification, detection, and segmentation, among others, whereas generative models additionally demand the capability to produce novel content grounded in perceptual understanding.
@@
Compared to traditional discriminative models that focus on perceptual tasks, generative models in continual learning settings are expected to go a step further, as shown in Figure 1 (b), by creating new content grounded in understanding and recognition. We summarize their increased challenges in continual learning into two main aspects: (1) More complex modeling objectives. Discriminative methods focus on modeling posterior and discriminative probabilities (i.e., $P(y|x)$), where catastrophic forgetting primarily manifests as a drift in the decision boundary within the classification space (Zhu et al., 2025; Zhou et al., 2024a; Liu et al., 2025b; Zhang et al., 2023c; Guo et al., 2024c; Zhu et al., 2021a). In contrast, generative models aim to grasp the underlying probability distribution (i.e., $P(x)$), which requires continual learning to maintain both semantic coherence and knowledge completeness within an open-ended generative space. To achieve this, it is essential to refine the generative reasoning process alongside the underlying factual knowledge representation, thereby mitigating hallucinations (Huang et al., 2025c; Bai et al., 2024) and semantic inconsistencies (Zhang et al., 2024a). (2) Greater diversity in task formats. Generative models encounter a wide variety of tasks during continual learning, with differences arising from factors such as new input modalities, novel domain knowledge, and evolving task characteristics. This variability increases the risk of catastrophic forgetting and places greater demands on multi-task knowledge integration. In contrast, discriminative models typically focus on a narrow range of tasks, such as classification, detection, or segmentation, which makes continual learning comparatively less challenging. Figure 2 demonstrates the continual learning scenarios across four representative generative models. Despite their impressive capabilities, enabling these models to learn continuously remains challenging and underexplored.
Several surveys focus on addressing continual learning and catastrophic forgetting, but they exhibit certain limitations. First, reviews focusing on discriminative models have attracted significant interest (Masana et al., 2022; Wang et al., 2024c; Zhou et al., 2024b; Menezes et al., 2023; Yuan and Zhao, 2024), but due to fundamental differences between discriminative and generative paradigms, these methods cannot be straightforwardly applied to continual learning in generative models. Furthermore, existing surveys on continual learning for generative models generally concentrate on specific architectures, such as LLMs (Wu et al., 2024b; Shi et al., 2024; Zheng et al., 2025d; Yang et al., 2025) or MLLMs (Yu et al., 2024; Huo and Tang, 2025; Huang et al., 2025b). Although these works provide systematic overviews within their respective domains, they often lack comprehensive discussions on the relationships, similarities, and differences among various types of generative models. To fill this gap, in this work, we present the first comprehensive review of continual learning across a wide spectrum of mainstream generative models, including LLM, MLLM, VLA, and Diffusion Model. Specifically, we start by uncovering the underlying principle in the context of continual learning shared across these types of models from a unified perspective. Subsequently, we systematically analyze the key approaches and challenges in each type of generative model, respectively. As a whole, our paper provides up-to-date advances in the literature and offers a holistic perspective for future works.
@@
Figure 2. Illustrations of continual learning for generative AI models: Adapting to real-world demands by learning new instructions, knowledge, modalities, actions, and creations without catastrophic forgetting.
@@
The paper is organized as follows: In Section 2, we present the setup of continual learning, covering its basic formulation, evaluation metrics, and taxonomy. Sections 3 through 6 provide an in-depth discussion of continual learning in Large Language Models (Section 3), Multimodal Large Language Models (Section 4), Vision-Language-Action Models (Section 5), and Diffusion Models (Section 6). Each domain is examined in terms of training strategies for continual learning tasks, representative benchmarks, model architectures, as well as the motivations and practical implementations of existing approaches. In Section 7, we further explore potential future research directions and emerging trends in continual learning methods built upon generative models. Finally, Section 8 concludes the paper and introduces a continuously updated GitHub repository for tracking the latest progress in continual learning for generative models.
2. Overview of Continual Learning
2.1. Basic Formulation
Continual learning tackles the challenge of acquiring knowledge from dynamically evolving data distributions, where training samples from distinct tasks arrive sequentially. A continual learning model is required to assimilate new knowledge while preserving previously learned information, typically under constraints that limit or prevent access to past data, while maintaining strong performance across all tasks. Formally, suppose there are $T$ continual learning tasks in total. The training samples for task $t$ can be represented as $\mathcal{D}_{t}=\{(x,y)\}$, where $x$ is the input data and $y$ is the corresponding label. Notably, the types of training data required vary across different generative models. For example, in MLLM, $(x,y)$ typically consists of a triplet comprising input images, user instructions, and target answers $(\textbf{X}_{v},\textbf{X}_{ins},\textbf{X}_{a})$. In VLA models, it can take the form of state-action pairs $(\mathcal{S},\mathcal{A})$, where the state $\mathcal{S}$ includes an image observation and a text instruction. During the training of task $t$, only the data for the current task is accessible. The training objective of each continual learning task can be formalized as follows:
(1) $\mathop{\arg\min}_{\theta}~\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))\right],$
2.2. Evaluation Metrics
The evaluation of a model’s continual learning ability typically involves three aspects: overall performance on all learned tasks, the extent of forgetting on previously learned tasks, and the generalization capability to unseen tasks. For clarity, we denote $a_{t,k}$ as the performance on the $k$-th task after the model has been sequentially trained from task 1 to task $t$ (Wang et al., 2024c).
(2) $\text{Avg}_{t}=\frac{1}{t}\sum_{j=1}^{t}\text{Last}_{j},~\text{Last}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$
(3) $f_{k,t}=\max_{i\in\{1,\cdots,t-1\}}(a_{i,k}-a_{t,k}),\forall k<t.$
(4) $\text{FM}_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}f_{i,t}.$
(5) $\text{BWT}_{t}=\frac{1}{t-1}\sum_{k=1}^{t-1}(a_{t,k}-a_{k,k}).$
FM quantifies catastrophic forgetting, where lower values indicate better retention of previous knowledge. BWT evaluates knowledge transfer between tasks, with positive values (i.e. BWT > 0) reflecting beneficial forward transfer to prior tasks and negative values (i.e. BWT < 0) indicating interference.
Generalization capability can be evaluated by zero-shot transfer (ZT), which is designed to quantify how well zero-shot transferability is maintained throughout the training process. Formally, for task $k$, it calculates the average accuracy $A_{k}$ of models trained on preceding tasks evaluated on task $k$:
(6) $A_{k}=\frac{1}{k-1}\sum_{i=1}^{k-1}a_{i,k},\forall k>1.$
ZT is the average of $A_{k}$ after training on the final task:
(7) $\text{ZT}=\frac{1}{T-1}\sum_{k=2}^{T}A_{k}.$
It is worth noting that the calculation of $a_{t,k}$ depends on the specific task type. For example, Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) for text generation task, Intersection-over-Union (IoU) (Cermelli et al., 2020) for visual grounding task, Fréchet Inception Distance (FID) (Heusel et al., 2017) for text-to-image generation task, and so on. There are several other evaluation criteria available, and we refer readers to their original papers for detailed specifications.
@@ 2.2. Evaluation Metrics
(§ 3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.5cm [Architecture-based
@@
(§ 3.2), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.5cm [TPEM (Geng et al., 2021), Zheng et al (Zheng et al., 2025b), TGL (Fernandez et al., 2024), LOIRE (Han et al., 2025), AdapterCL (Madotto et al., 2020), SAPT (Zhao et al., 2024b), MoRAL (Yang et al., 2024a), I-LoRA (Ren et al., 2024), SLIM (Han et al., 2024), SEE (Wang et al., 2025a),
@@
P-Prompts (Razdaibiedina et al., 2023), Q-Tuning (Guo et al., 2024b), WAVE-CRE (Le et al., 2025) , color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.25cm] ] [Regularization-based
@@
(§ 3.3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.5cm [O-LoRA (Wang et al., 2023a), DAS (Ke et al., 2023), ARPER (Mi et al., 2020), SEEKR (He et al., 2024a), TaSL (Feng et al., 2024), FVG (Jiang et al., 2025),
@@
DaCL (Wu et al., 2023a), DYNAINST (Mok et al., 2023), Velocitune (Luo et al., 2024), Recurrent-KIF (Feng et al., 2025), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.25cm] ] [Replay-based
@@
(§ 3.4), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.5cm [CT0 (Scialom et al., 2022), Huang et al. (Huang et al., 2024c), Anh et al. (Anh et al., 2025), InsCL (Wang et al., 2024b), HESIT (Chen et al., 2024d),
@@
Sun et al. (Sun and Gao, 2024), D-CPT Law (Que et al., 2024), LAMOL (Sun et al., 2019), HMI-LAMOL (Maekawa et al., 2023), PCLL (Zhao et al., 2022), DCL (Zeng et al., 2023), SSR (Huang et al., 2024a), SynE (Chen et al., 2024a), KPIG (He et al., 2024b), Guo et al. (Guo et al., 2024a) , color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.25cm] ] ] [Multimodal Large Language Models
@@
(§ 4), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.5cm [Architecture-based
@@
(§ 4.2), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.5cm [Eproj (He et al., 2023), TAM-CL (Cai et al., 2023), MoELoRA (Chen et al., 2024f), Continual LLaVA (Cao et al., 2024), HiDe-LLaVA (Guo et al., 2025b), LLaCA (Qiao et al., 2024), SMoLoRA (Wang et al., 2024a), CL-MoE (Huai et al., 2025),
@@
LLaVA-CMoE (Zhao et al., 2025a), DISCO (Guo et al., 2025c), TRIPLET (Qian et al., 2023), ModalPrompt (Zeng et al., 2024b),
@@
Fwd-Prompt (Zheng et al., 2024), ColPro (Cai et al., 2024), CluMo (Cai and Rostami, 2024), MR-LoRA (Zhao et al., 2025b), D-MoLE (Ge et al., 2025), BranchLoRA (Zhang et al., 2025b), ProgLoRA (Yu et al., 2025a) , color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.25cm] ] [Regularization-based
@@
(§ 4.3), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.5cm [SCD (Lao et al., 2023), CS-VQLA (Bai et al., 2023b), Chen et al. (Chen et al., 2024b), CrossSDC (Pian et al., 2024b), MAFED (Nikandrou et al., 2024),
MoInCL (Pian et al., 2024a), QUAD (Marouf et al., 2025), Model Tailor (Zhu et al., 2024c), SPIDER (Huang et al., 2024b), SEFE (Chen et al., 2025a),
@@
LoRASculpt (Liang et al., 2025), LLaVA-c (Liu et al., 2025c) , color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.25cm] ] [Replay-based
@@
(§ 4.4), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.5cm [VQACL (Zhang et al., 2023d), SGP (Lei et al., 2023), ProtoGroup (Zhang et al., 2025a), Lin et al. (Lin et al., 2025), Adapt-$\infty$ (Maharana et al., 2024),
@@
OASIS (Lee et al., 2025), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.25cm] ] ] [Vision-Language-Action
Models
@@
(§ 5), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.5cm [Architecture-based
@@
(§ 5.2), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.5cm [LIBERO (Liu et al., 2023b), Lotus (Wan et al., 2024), QueST (Mete et al., 2024), LEGION (Meng et al., 2025), Jia et al. (Jia et al., 2025),
@@
DRAE (Long et al., 2025) , color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.25cm] ] [Regularization-based